# Retrieve to Explain: Evidence-driven Predictions with Language Models

### v2

@@ 4.5.3 Results
R2E-uncor Genetics 0.579
R2E-cor Combined 0.633
R2E-audited-cor Combined 0.638
Primary results are shown in Table 3, with further results across temporal training, bias correction, and method of combining modalities available in Appendix T. Previous analyses restrict clinical trial outcomes to those diseases with genetics data (Minikel et al., 2023); as literature evidence is not restricted to those diseases, we instead show performance against all diseases in the Clinical Trial Outcomes data. Nonetheless, results are qualitatively similar on the restricted set (Table S9).
When reasoning from templated genetics evidence alone, we found that R2E displays improved performance over the genetics baseline despite leveraging the same data. However, R2E had much stronger performance when using the biomedical literature. There was only a marginal improvement from combining templated genetics evidence and the biomedical literature (see additional approaches to combining data sources, with similar performance, in Table S10). There was variance in relative performance across disease areas and modalities (Figure S4).
Shapley values were correlated with GPT-4 relevance annotations, albeit less strongly than in Section 4.4.2, with an AUROC of 0.665. GPT-4 had an accuracy of 82.2% compared to human expert annotations.
4.6 Auditing Explanation Evidence with GPT-4
4.6.1 Dataset Construction
We sought to assess the hypothesis that R2E explanations could enable human- or LLM-in-the-loop feedback to remove false positives. Pooling R2E predictions on the Clinical Trial Outcomes dataset, we used GPT-4 to annotate the relevance of 20,000 query-evidence pairs with the highest Shapley values (computed on pre-sigmoid outputs). We then reran R2E inference on the full dataset, replacing evidence labelled as irrelevant with the NULL embedding. See Appendices Q and U for the GPT-4 prompt and examples.
4.6.2 Results
Removing irrelevant sentences from the coalitions used for R2E predictions resulted in a small but significant increase in AUROC (Table 3, de Long test, $p=0.004$).
In addition, 527/809 high Shapley value genetics sentences were annotated as relevant, and 268/527 of these related to a disease that was neither a substring of, nor contained, the Clinical Trial Outcomes disease.
5 Limitations and Future Work
Retrieving evidence at inference time to make predictions has a cost: generating a score for each answer requires a vector search over the answer’s evidence, followed by a model forward pass. In comparison, predicting from a multiclass model (like the MLM) requires a single forward pass without retrieval. However, retrieval and reasoning are both independent across possible answers, so R2E scales effectively with computational resources.
Retrieval-based inference also has flexibility benefits beyond those explored here. By filtering retrieved evidence on document metadata, users could customize the ranking at inference time; with a scientific literature dataset, this could include filtering supporting evidence to specific timespans, publications, impact factors, paper sections, or keywords.
In Sections 4.4 and 4.5, we applied the model directly to downstream tasks; in the case of clinical trials, we simply adopted a one-size-fits-all query template. Instead, the system could be fine-tuned for the task of interest. Fine-tuning is of particular interest in a human-in-the-loop setting, where a user could identify faulty evidence use (as opposed to a faulty prediction) and tailor the feedback accordingly. Similarly, an LLM could be used to generate labels for model fine-tuning in addition to inference-time auditing.
The evidence templating approach used for genetics is relatively general, and could be applied to other data modalities (e.g. transcriptomics evidence in drug discovery). However, care should be taken with respect to the distribution of the training data (e.g. in drug discovery, templating evidence consistently with how it might be discussed in the literature).
Finally, given the small size of the Retriever and Reasoner architectures (10 and 2 million parameters respectively), we expect improvements with model scaling. Similarly, longer literature passages (e.g. paragraphs rather than sentences) are likely to provide more context and improve performance.
6 Discussion
By retrieving evidence to make predictions, R2E enables explainability, a critical feature in complex, high-risk decision-making. R2E also outperformed the target identification baselines after accounting for class imbalance. Combined with the proposed bias correction technique, this improves the ability to make informed predictions about novel and understudied, but promising targets. Finally, R2E outperformed competing approaches on the difficult task of predicting clinical trial efficacy outcomes, without task-specific fine-tuning. Performance was further improved by auditing R2E’s prediction explanations using GPT-4, an approach made possible by the retrieval-based setup.
@@ 6 Discussion
When R2E reasoned from genetics data templated into natural language, we found that it significantly outperformed existing approaches (Nelson et al., 2015) leveraging the same data. This could be explained by the language model’s capacity to leverage “soft” semantic associations (e.g. recognizing correlations between diseases beyond those identified by ontological similarity of diseases), as corroborated by the inspection of high-scoring genetics evidence (see Appendix T and Figure S5 for further analysis). This is particularly important given the relatively poor performance of the genetics dataset across all clinical outcomes; while direct genetics evidence is predictive when it exists, it often does not exist or is not applicable. In contrast, the scientific literature has broad coverage across a wide range of indications.
The recent explosion in retrieval-based approaches has challenged the paradigm of using a single model to both memorize factual knowledge and reason from that knowledge. We show here that retrieval can convey not only performance and flexibility advantages, but also significantly improved transparency into how the model reasons from evidence.
Impact Statement
In general, this paper presents work whose goal is to advance the field of Machine Learning. There are many potential widely discussed societal consequences of developments in machine learning, that we do not think warrant specifically highlighting here.
@@
In this paper, we demonstrate the application of R2E to predict and explain protein-coding genes in response to a user query. This application is quite different to either the generality of large language models or the structural biology and chemistry foci of the AI-enabled biological tools most typically associated with any potential dual risk concern. As with other tools that facilitate biomedical research and understanding, the ability to identify and understand particular genes could be applied in a range of use cases. For this paper, we do not believe there to be material risks to highlight, especially noting: (1) We are not releasing proprietary training data, code, or model weights; (2) Explanations provided by R2E are either publicly-available extracts from the scientific literature or structured data templated in natural language, and can be interpreted by expert users in the context of their wider biomedical understanding, but do not significantly lower the barrier to entry for non-experts users; (3) R2E is predicting at the level of drug targets, with multiple complex downstream steps required to translate the identification of a target that may achieve a particular biological effect, into a capability to intervene on that target.
Principles for the ethical deployment of AI at BenevolentAI are outlined on our website.
Acknowledgements
@@
The authors would like to thank Nicola Richmond and Julien Fauqueur for their helpful comments and feedback during drafting, Bradleigh Whitton and James Grey for applying their biological expertise to annotate evidence for validation of GPT-4 Evidence Annotations datasets, Alison McGarvey and Eryk Kropiwnicki for their help and advice on the process for templating genetics data into sentences, Antonios Poulakakis Daktylidis for their assistance with validation of the Clinical Trial Outcomes data, and Hao-Chih Lee for their helpful insights on evaluations of evidence explanations.
@@ Appendix A Accessing Evaluation Datasets
  • Gene Description Facts © 2024 by BenevolentAI is licensed under CC BY-NC-SA 4.0. It is adapted from ”Universal Protein Resource (UniProt)” by Uniprot Consortium, used under CC BY 4.0.
@@ Appendix B Masked Entity-Linked Corpus, Dataset Splits & Sizes
We created three distinct corpus splits $\mathcal{S}_{1}$, $\mathcal{S}_{2}$, and $\mathcal{S}_{3}$ to handle training, validation and testing across the both MLM/Retriever and Reasoner (Figure S3). These splits were generated at the level of entire documents to reduce the occurrence of highly similar sentences between splits.
For the MLM/Retriever, datasets were composed as follows:
@@
For R2E, we use $\mathcal{D}^{e}$ to refer to an evidence/retrieval corpus and $\mathcal{D}^{q}$ to refer to a query corpus. The datasets were composed as follows:
@@
  • $\mathcal{D}^{q}_{test}\subset\mathcal{S}_{3}:\mathcal{D}^{q}_{test}\cap\mathcal{D}^{q}_{val}=\emptyset$
$\mathcal{S}_{3}$ was partitioned from $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, either randomly (with 1.5 million sentences allocated to $\mathcal{S}_{3}$) or using a year split (all documents from a specified year onwards allocated to $\mathcal{S}_{3}$ and all documents from prior years allocated to $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$). $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ were always created by a random partition with 1.5 million sentences allocated to $\mathcal{S}_{2}$ and the remainder in $\mathcal{S}_{1}$.
The number of sentences in $\mathcal{S}_{1}$ for each of the year splits used in our reported experiments, corresponding to particular evaluations, were:
  • 2005 year split (${\mathcal{D}}^{q}_{eval}=$ Clinical Trial Outcomes): 16.2 million
  • 2020 year split (${\mathcal{D}}^{q}_{eval}={\mathcal{D}}^{q}_{test}=$ Held-out Biomedical Literature): 112 million
  • Full data, no year split (${\mathcal{D}}^{q}_{eval}=$ Gene Description Facts): 157 million
The above splitting procedure is illustrated in Figure S3 for the case of a 2020 year split model. In the 2020 year split case, note the disjoint subsets sampled from $\mathcal{S}_{3}$ and used to create overall validation ($\mathcal{D}^{q}_{val}$) and test ($\mathcal{D}^{q}_{test}$) queries, for ranking metric evaluations over all genes in Genes; namely the:
  • Held-out Biomedical Literature validation dataset: 1 sentence per gene, sampled from publicly-available abstract section sentences from 2020 onwards. Used for hyperparameter selection and ablation experiments described in Appendices D & J respectively.
  • Held-out Biomedical Literature test dataset: 1 sentence per gene per year for 2020 onwards, sampled from publicly-available abstract section sentences. Used for evaluations described in Section 4.3.1, including evaluation of the MLM and other baselines.
Figure S3: Masked entity-linked corpus data splits with a 2020 split year. Here we illustrate how the masked entity-linked corpus was partitioned to enable MLM/Reasoner and Retriever training, validation, and testing. Specifically the example of a 2020 year split dataset is shown, with overall validation and testing of R2E (and all baselines) using queries from 2020 onwards ($\mathcal{D}^{q}_{val}$ and $\mathcal{D}^{q}_{test}$, disjoint subsets sampled from Split 3). The overall validation and test query sets here correspond to the Held-out Biomedical Literature validation and test datasets. The partitioning setup shown here for the 2020 year split, is illustrative of other data split setups used in our experiments - the key differences being that the Clinical Trial Outcomes evaluation used a 2005 year split dataset, and Gene Description Facts evaluation used a random split rather than a year split dataset.
@@ Appendix D R2E Hyperparameters
The R2E model was implemented using PyTorch deep learning library (Paszke et al., 2019).
@@
All sentences were tokenized, and then truncated and padded to a length of 128, using the same vocabulary as PubMedBERT (Gu et al., 2021). Pre-processing of training examples for both Retriever and Reasoner training was done using Apache Spark (Zaharia et al., 2016). Both the Retriever and Reasoner were trained independently, for 10 epochs on a single Tesla V100 GPU, with a total training time of approximately 1 week.
@@
The final R2E Retriever architecture, as well as the MLM baseline, consisted of a scaled down version of PubMedBERT (Gu et al., 2021) trained from scratch on the task described in 3.2, with 4 hidden layers, 4 attention heads, an intermediate size of 512, a hidden size of 256, and total size of 10 million parameters. Final Retriever/MLM training used a batch size of 512, a categorical cross-entropy loss, and an AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of 0.0001 and no weight decay.
@@
The R2E architecture is summarised in Figure 2. The final query-evidence encoder component of the R2E Reasoner architecture consisted of a layer normalisation across all concatenated query/evidence pairs, then two conv1d layers each with kernel size of 1 (first layer: 2 input channels, 8 output channels; second layer: 8 input channels, 1 output channel) across each query/evidence pair individually. The final evidence combiner component of the R2E Reasoner architecture consisted of a set transformer (Lee et al., 2019) over all query-evidence embeddings returning a single embedding, followed by a linear layer and sigmoid to output a binary probability. The set transformer had 4 heads, 2 induced set attention blocks with 32 inducing points for the encoder, and a pooling by multihead attention followed by two set attention blocks in the decoder. The Reasoner had a total size of 2 million parameters. After freezing the Retriever weights, the final Reasoner training used a batch size of 2048, binary cross-entropy loss, and AdamW optimizer with a learning rate of 0.0001 and weight decay of 0.001. For both training and inference, 64 evidence sentences were retrieved for a given query. A post-hoc frequency bias correction factor of 0.5 was used for the R2E-Corrected variant (Section 3.6 and Appendix G for details of post-hoc correction).
@@
The post-hoc frequency bias correction factor selection and architectural comparison ablations (Appendix J) were based on MRR for a 2020 year split model, on a Held-out Biomedical Literature validation set containing one biomedical literature cloze-style query sentence per gene in Genes from publicly-available abstract sections (see Appendix B). The resulting 15477 validation set queries were therefore sentences published from 2020 onwards, and retrieval corpus sentences published prior to 2020. The learning rate was chosen to reduce training time while maintaining training stability, and the batch size selected to optimise GPU utilisation. We did not evaluate variations of model scale and leave this to future work.

### v4

@@ 4.5 Clinical Trial Outcomes
R2E-cor Both 0.633
R2E-audit Both 0.638
Results
Table 2 shows primary results, while Appendix U includes several further results and detailed discussions, including on relative success (Appendices U.1-U.3; Figure 4). Overall, R2E variants incorporating biomedical literature matched or outperformed all baselines.
@@ Results
Notably, R2E significantly outperformed the widely-used genetics baseline (Genetic) when leveraging only the exact same underlying genetics data templated into sentences (R2E-uncor (genetic); $p<0.001$). This could be explained by the language model’s capacity to leverage “soft" semantic associations (e.g. recognizing correlations between diseases / traits beyond ontological similarity), as corroborated by the inspection of high-scoring genetics evidence (Appendix U.7; Figure 6). The addition of literature resulted in a significant further improvement ($p<0.001$). The relative under-performance of models using genetics data alone compared to those using biomedical literature likely reflects the lack of genetic coverage of diseases, despite it being predictive when available. In contrast, the literature has broad coverage across diseases. Figure 5 (Appendix U.5) shows performance by disease area with greater variability for genetics.
R2E also significantly outperformed the few-shot, chain-of-thought prompted GPT-4 baseline with retrieval augmentation. The full method and results for this baseline are described in Appendix V and U.4 respectively.
There was only a marginal improvement from combining templated genetics evidence and the biomedical literature over literature alone. This could be explained by the $\sim$200:1 balance of literature to genetics-derived sentences in the evidence corpus, and the potential redundancy of the genetics evidence given information already represented in the literature. Additional approaches to combining data sources, with similar performance, are compared in Appendix U.6 (Table 9).
Evidence Shapley values correlated with binary GPT-4 relevance annotations (AUROC: 0.665) and GPT-4 with human-expert annotations ($82.2$% agreement rate). Together, this suggests moderate agreement on evidence relevance. See Appendix S for a case study of examples.
4.6 Auditing Explanation Evidence
We sought to assess the hypothesis that R2E explanations could enable human- or LLM-in-the-loop feedback to remove false positive evidence. Pooling R2E predictions on the Clinical Trial Outcomes dataset, we used GPT-4 to annotate the relevance of 20,000 query-evidence pairs with the highest Shapley values (computed on pre-sigmoid outputs). We then reran R2E-cor inference on the full dataset, replacing evidence labelled as irrelevant with the NULL embedding, yielding a small but significant improvement (R2E-audit, Table 2, $p=0.004$). Said differently, by allowing evidence to be audited, R2E’s explainability enabled further performance improvement. For the GPT-4 prompt, and auditing examples, see Appendices R and W.
5 Conclusions
By retrieving evidence to make predictions, R2E enables faithful and quantitative explainability, a critical feature in complex, high-stakes settings such as drug target identification. R2E matched or outperformed all baselines across the three target identification evaluation tasks. Combined with the proposed bias correction technique, this improves the ability to make informed predictions about novel and understudied, but promising targets. Finally, R2E outperformed a widely-used competing approach on the important and challenging task of predicting clinical trial efficacy outcomes, without task-specific fine-tuning. Performance was further improved by auditing R2E explanations using GPT-4, an approach made possible by the retrieval-based setup. We show here that retrieval can provide not only performance and flexibility advantages, but also significantly improved transparency into how the model reasons from evidence.
6 Limitations
Retrieving evidence at inference time to make predictions has a cost: each answer score requires a vector search over the answer’s evidence, followed by a model forward pass. In comparison, predicting with a multiclass model (MLM) requires a single forward pass without retrieval. For efficient scaling, retrieval and reasoning can be parallelized across answers (Appendix H).
Retrieval-based inference has flexibility benefits beyond those explored here. By filtering retrieved evidence on document metadata, users could customize the ranking at inference time; with a scientific literature dataset, this could include filtering supporting evidence to specific timespans, publications, impact factors, paper sections, or keywords.
The performance of a retrieval-based approach is expected to be sensitive to the completeness of the underlying corpus. However, R2E explanations help to make limitations or biases of the corpus more transparent than would be the case for a fully parametric approach, and parametric approaches are also sensitive to their training corpus.
In Sections 4.4 and 4.5, we applied the model directly to downstream tasks; in the case of clinical trials, we simply adopted a one-size-fits-all query template. Instead, the system could be fine-tuned for the task of interest. Fine-tuning with human feedback is of particular interest here, since with R2E a user can focus on faulty evidence use (as opposed to a faulty prediction). Similarly, an LLM could be used to generate evidence-level labels for model fine-tuning in addition to the inference-time auditing described in Section 4.6.
The evidence templating approach used for genetics in Section 4.5 is relatively general, and could be applied to other data modalities, such as transcriptomics evidence in drug discovery. We focused on genetics because it is well-established as being predictive of clinical trial outcomes. For new modalities, care should be taken with respect to the distribution of the training data. For example, for scientific applications, evidence should be templated consistently with how it might be discussed in the literature corpus.
Performance gains might be made by scaling the Retriever and Reasoner, as well as extending to longer literature passages to increase context, for example paragraphs instead of sentences.
7 Ethical Considerations
As detailed in Section 1, the explainability of R2E has the potential to positively impact the utility and adoption of models in high-stakes human-in-the-loop settings where explainability is often paramount, as exemplified by target identification. For target identification specifically, the improvements here could have significant positive consequences for the success of drug development programs and therefore the rate at which new more efficacious therapies become available to patients.
@@ 7 Ethical Considerations
The application of R2E to predict and explain protein-coding genes in response to a user query is quite different to either the generality of large language models or the structural biology and chemistry foci of the AI-enabled biological tools most typically associated with any potential dual risk concern. As with other tools that facilitate biomedical research and understanding, the ability to identify and understand particular genes could be applied in a range of use cases. For this paper, we do not believe there to be material risks to highlight, especially noting: (1) We are not releasing proprietary training data, code, or model weights; (2) Explanations provided by R2E are either publicly-available extracts from the scientific literature or non-textual data templated in natural language, and can be interpreted by expert users in the context of their wider biomedical understanding, but do not significantly lower the barrier to entry for non-experts users; (3) R2E is predicting at the level of drug targets, with multiple complex downstream steps required to translate the identification of a target that may achieve a particular biological effect, into a capability to intervene on that target.
Acknowledgments
@@
The authors would like to thank Nicola Richmond and Julien Fauqueur for their helpful comments and feedback during drafting, Bradleigh Whitton and James Grey for applying their biological expertise and consenting to the use of their evidence annotations for validation of GPT-4 Evidence Annotations datasets (Clinical Trial Outcomes and Gene Description Facts respectively), Alison McGarvey and Eryk Kropiwnicki for their help and advice on the process for templating genetics data into sentences, Antonios Poulakakis Daktylidis for their assistance with validation of the Clinical Trial Outcomes data, and Hao-Chih Lee for their helpful insights on evaluations of evidence explanations.
@@ Appendix A Accessing Evaluation Datasets
We make the three performance evaluation datasets used in this paper publicly available as part of the Supplementary Material, licensed under CC BY-NC-SA 4.0. Specific licensing information for the datasets is as follows:
  • Clinical Trials Outcomes is licensed under CC BY-NC-SA 4.0. We have permission from Citeline PharmaProjects to publicly release the subset of their data that is used here.
  • Gene Description Facts is licensed under CC BY-NC-SA 4.0. It is adapted from "Universal Protein Resource (UniProt)" by Uniprot Consortium, used under CC BY 4.0.
@@
  • Gene Description Facts © 2024 by BenevolentAI is licensed under CC BY-NC-SA 4.0. It is adapted from "Universal Protein Resource (UniProt)" by Uniprot Consortium, used under CC BY 4.0.
@@ Appendix B Masked Entity-Linked Corpus, Dataset Splits & Sizes
We created three distinct corpus splits $\mathcal{S}_{1}$, $\mathcal{S}_{2}$, and $\mathcal{S}_{3}$ (Figure 3). These splits were generated at the level of entire documents to reduce the occurrence of highly similar sentences between splits.
For Held-out Biomedical Literature (Appendix J) and Clinical Trial Outcomes (Section 4.5) experiments, where evaluation queries were associated with metadata for year of publication and earliest clinical development date respectively, a temporal year split setup was used to ensure models trained on and retrieved from sentences prior to the start year of the evaluation data. Specifically, for these year split experiments, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ were random samples from before the split year with 1.5 million sentences allocated to $\mathcal{S}_{2}$ and the remainder to $\mathcal{S}_{1}$. $\mathcal{S}_{3}$ contained all sentences from documents after the split year. A split year of 2005 was used for Clinical Trial Outcomes ($|\mathcal{S}_{1}|=$ 16.2 million sentences), and a split year of 2020 for Held-out Biomedical Literature ($|\mathcal{S}_{1}|=$ 112 million sentences).
For Gene Description Facts experiments (Section 4.4), where evaluation queries did not correspond to a particular year, no year split was used. Specifically, $\mathcal{S}_{1}$, $\mathcal{S}_{2}$, and $\mathcal{S}_{3}$ were all random samples of the corpus, with 1.5 million sentences allocated to each of $\mathcal{S}_{2}$ and $\mathcal{S}_{3}$, and the remainder to $\mathcal{S}_{1}$ (157 million sentences).
Training, validation and testing datasets were then constructed for both R2E Retriever / MLM and R2E Reasoner, by using the appropriate $\mathcal{S}_{1}$, $\mathcal{S}_{2}$, and $\mathcal{S}_{3}$ splits.
For the R2E Retriever / MLM, training and validation datasets were composed as follows:
@@
For the R2E Reasoner, for each of train, validation and test, both retrieval and query corpora were needed, to ensure query sentences were not also included in the retrieval corpus. We use $\mathcal{D}^{e}$ to refer to a retrieval corpus of evidence sentences and $\mathcal{D}^{q}$ to refer to the query corpus of sentences. The datasets were composed as follows:
@@
  • $\mathcal{D}^{q}_{test}\subset\mathcal{S}_{3}:\mathcal{D}^{q}_{test}\cap\mathcal{D}^{q}_{val}=\emptyset$, i.e. a held-out subset of $\mathcal{S}_{3}$, without overlap with $\mathcal{D}^{q}_{val}$
The above splitting procedure is illustrated in Figure 3 for the case of the 2020 year split setup used for Held-out Biomedical Literature experiments. For this Held-out Biomedical Literature setup, the disjoint subsets sampled from $\mathcal{S}_{3}$ and used to create overall validation ($\mathcal{D}^{q}_{val}$) and test ($\mathcal{D}^{q}_{test}$) queries, are those used to report ranking metric evaluations over all genes in Genes; namely the:
  • Held-out Biomedical Literature validation dataset: 1 sentence per gene, sampled from publicly-available abstract section sentences from 2020 onwards. Used for hyperparameter selection and ablation experiments described in Appendices D & L respectively.
  • Held-out Biomedical Literature test dataset: 1 sentence per gene per year for 2020 onwards, sampled from publicly-available abstract section sentences. Used for evaluations described in Section 4 and Appendix J, including evaluation of the MLM and other baselines.
Note the key difference between this 2020 year split setup for Held-out Biomedical Literature, and the setups for the other two evaluation datasets were:
  • Different year splits (as described above)
  • The queries used in evaluation were derived from those specific evaluation datasets, not a held-out split of the literature corpus (i.e. ${\mathcal{D}}^{q}_{eval}\neq{\mathcal{D}}^{q}_{test}$)
Figure 3: Masked entity-linked corpus for Held-out Biomedical Literature experiments. Here we illustrate how the masked entity-linked corpus was partitioned to enable Reasoner/MLM and Retriever training, validation, and testing. Specifically the example of a 2020 year split setup is shown, as was used for Held-out Biomedical Literature experiments.
@@ Appendix D R2E Hyperparameters
The R2E model was implemented using PyTorch deep learning library Paszke et al. (2019).
@@
All sentences were tokenized, and then truncated and padded to a length of 128, using the same vocabulary as PubMedBERT Gu et al. (2021). Pre-processing of training examples for both Retriever and Reasoner training was done using Apache Spark Zaharia et al. (2016). The Retriever and Reasoner were trained sequentially, each for 10 epochs on a single Tesla V100 GPU, with a total training time of approximately 1 week.
@@
The final R2E Retriever architecture, as well as the MLM baseline, consisted of a scaled down version of PubMedBERT Gu et al. (2021) trained from scratch on the task described in 3.2, with 4 hidden layers, 4 attention heads, an intermediate size of 512, a hidden size of 256, and total size of 10 million parameters. Final Retriever/MLM training used a batch size of 512, a categorical cross-entropy loss, and an AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of 0.0001 and no weight decay.
@@
The R2E architecture is summarised in Figure 2. The final query-evidence encoder component of the R2E Reasoner architecture consisted of a layer normalisation across all concatenated query/evidence pairs, then two conv1d layers each with kernel size of 1 (first layer: 2 input channels, 8 output channels; second layer: 8 input channels, 1 output channel) across each query/evidence pair individually. The final evidence combiner component of the R2E Reasoner architecture consisted of a set transformer Lee et al. (2019) over all query-evidence embeddings returning a single embedding, followed by a linear layer and sigmoid to output a binary probability. The set transformer had 4 heads, 2 induced set attention blocks with 32 inducing points for the encoder, and a pooling by multihead attention followed by two set attention blocks in the decoder. The Reasoner had a total size of 2 million parameters. After freezing the Retriever weights, the final Reasoner training used a batch size of 2048, binary cross-entropy loss, and AdamW optimizer with a learning rate of 0.0001 and weight decay of 0.001. For both training and inference, 64 evidence sentences were retrieved for a given query. A post-hoc frequency bias correction factor of 0.5 was used for the R2E-cor variant (Section 3.6 and Appendix G for details of post-hoc correction).
Table 3: R2E hyperparameter summary
Component Hyperparameters
Tokenization Max sequence length: 128 Truncated and padded Tokenized using PubMedBERT vocabulary
Training Setup 10 epochs for Retriever 10 epochs for Reasoner Single Tesla V100 GPU Total training time: approx. 1 week
Retriever / MLM Baseline Architecture: 4 hidden layers, 4 attention heads Hidden size: 256, Intermediate size: 512 Total parameters: approx. 10M Training: Batch size: 512 samples Loss: Categorical cross-entropy Optimizer: AdamW Learning rate: 0.0001, Weight decay: 0.0
Reasoner Encoder: LayerNorm over concatenated query-evidence pairs Two Conv1D layers: * 1st: 2 $\rightarrow$ 8 channels * 2nd: 8 $\rightarrow$ 1 channel Combiner (Set Transformer): 4 heads 2 ISABs (encoder) with 32 inducing points PMA pooling 2 SABs (decoder) Output: Linear layer + sigmoid Total parameters: approx. 2M Training: Batch size: 2048 samples Loss: Binary cross-entropy Optimizer: AdamW Learning rate: 0.0001, Weight decay: 0.001
Retriever Inference 64 evidence sentences retrieved per query
Bias Correction Post-hoc frequency bias correction factor: 0.5 for R2E-cor variant 0.0 for R2E-uncor variant
@@
The post-hoc frequency bias correction factor selection and architectural comparison ablations (Appendix L) were based on MRR for a 2020 year split model, on a Held-out Biomedical Literature validation set containing one biomedical literature cloze-style query sentence per gene in Genes from publicly-available abstract sections (Appendix B). The resulting 15477 validation set queries were therefore sentences published from 2020 onwards, and retrieval corpus sentences published prior to 2020. The learning rate was chosen to reduce training time while maintaining training stability, and the batch size selected to optimise GPU utilisation. We did not evaluate variations of model scale and leave this to future work.